# Synth-to-Real Unsupervised Domain Adaptation for Instance Segmentation

### v2

@@ Abstract
Unsupervised Domain Adaptation (UDA) aims to transfer knowledge learned from a labeled source domain to an unlabeled target domain. While UDA methods for synthetic to real-world domains (synth-to-real) show remarkable performance in tasks such as semantic segmentation and object detection, very few were proposed for the instance segmentation task. In this paper, we introduce UDA4Inst, a model of synth-to-real UDA for instance segmentation in autonomous driving. We propose a novel cross-domain bidirectional data mixing method at the instance level to fully leverage the data from both source and target domains. Rare-class balancing and category module training are also employed to further improve the performance. It is worth noting that we are the first to demonstrate results on two new synth-to-real instance segmentation benchmarks, with 39.0 mAP on UrbanSyn Cityscapes and 35.7 mAP on Synscapes Cityscapes. UDA4Inst also achieves the state-of-the-art result on SYNTHIA Cityscapes with 31.3 mAP, +15.6 higher than the latest approach. Our code will be released.
@@
Unsupervised Domain Adaptation Instance Segmentation Synth-to-Real1 Introduction
@@
Instance segmentation provides a detailed and fine-grained understanding of the objects in an image, which is critical for a variety of real-world applications, such as multi-object tracking [29, 48, 2], pose estimation [5, 1], motion prediction [63, 35]. Specifically, in autonomous driving, the segmentation of instance objects is useful as it helps to locate safety-critical objects (such as pedestrians or riders) with accurate boundaries, enabling more advanced scene analysis and decision-making in many scenarios [8, 58, 2, 65].
@@
However, in practice, training instance segmentation models requires access to a large amount of densely labeled data. Each instance needs to be accurately delineated with pixel-level masks, which is laborious and highly time-consuming. In addition, in order to ensure a proper model generalization, the data must contain a wide range of object appearances, poses, scales, and backgrounds. Therefore, the label acquisition and object variability needed for instance segmentation encounter a data-hungry bottleneck in the real-world domain. This challenge demands a more intelligent approach to enhance the efficiency of the labeling of instance segmentation data.
@@
As an alternative, synthetic data is increasingly being used for deep learning applications as it enables large-scale training data for models of vision tasks [45]. Synthetic data offers labels annotated in a more time- and resource-efficient way, as it is directly generated by simulators [11, 44] along with images. Crucially, it allows for the incorporation of diversity into the training dataset, thereby enhancing the model’s generalization capability. Moreover, synthetic data can be tailored to produce scenarios that are rare in the real world.
@@
Utilizing synthetic data to replace real-world data for model training is based on the premise that training and test samples are drawn from the same distribution. Unfortunately, the distribution of these two domains of data is different (i.e. a domain gap does exist). UDA methods are essential for mitigating the inherent domain gap between the synthetic source and the real-world target domains. In recent years, many studies have explored the integration of synthetic data and UDA methods to address a wide range of computer vision tasks including image classification [39, 12, 40, 47, 42], object detection [54, 14, 16, 46, 67, 43], and semantic segmentation [3, 24, 25, 28, 17, 20, 38]. Comparatively, there has been relatively little research on UDA for instance segmentation, which is reflected in only three works [70, 27, 52] in the last five years. Among them, although [27, 52] do provide results on instance segmentation, they were models designed for the panoptic segmentation task, and [70] was based on a suboptimal MaskR-CNN framework with adversarial learning.
@@
In this paper, we focus on designing a novel synth-to-real UDA model for Instance Segmentation (UDA4Inst) following an online self-training pattern. We employ the advanced universal segmentation model Mask2Former [6] to demonstrate the effectiveness of our proposed UDA method. Unlike the previous models [56, 26, 13, 30] that only perform a unidirectional cross-domain data mixing, i.e. either source-to-target (S2T) or target-to-source (T2S), UDA4Inst employs a new bidirectional cross-domain data mixing based on instance objects. We show that our strategy improves the mAP results over the unidirectional ones. Furthermore, to address the ubiquity of the class imbalance problem in the existing datasets, UDA4Inst is equipped with a novel algorithm that can online balance rare-class instances during cross-domain data mixing. In this way, we observe a clear improvement in the segmentation for rare-class instance objects, which contributes to a better generalization of the model. As the third methodological contribution, we explore category module training where multiple classes are categorized into several modules for separate training, and the final results consist of the output from each module. Last but not least, we demonstrate the effectiveness of UDA4Inst on three UDA instance segmentation benchmarks. To the best of our knowledge, we are the first to perform UDA instance segmentation with synthetic datasets of UrbanSyn [15] and Synscapes [66], reaching respectively 39.0 mAP and 35.7 mAP on Cityscapes. Our results on SYNTHIA-to-Cityscapes reach the state-of-the-art with 31.3 mAP (56.9 mAP50), +15.6 (+22.6) higher than the previous best record [52].
@@ 2.1 Instance Segmentation
Early mainstream deep learning models based on convolutional neural networks (CNNs) for instance segmentation, such as MaskR-CNN [21] and YOLACT [4] have achieved remarkable successes. In recent years, with the adoption of the self-attention mechanism in natural language processing [60] and computer vision [10, 71, 49], transformer-based models [7, 6, 34, 31] have emerged and surpassed CNN-based architectures in many tasks such as object detection [34], semantic segmentation [7, 6, 34, 31], and panoptic segmentation [7, 6, 34, 31]. Among them, Mask2Former [6] has attracted great attention as a universal architecture that obtains top results in three major image segmentation tasks (panoptic, instance, and semantic) on benchmarks. It introduces masked attention, which extracts localized features by constraining cross-attention within predicted mask regions. In this work, we leverage this advanced universal segmentation architecture to prompt the state-of-the-art results in instance segmentation with UDA.
@@ 2.2 Unsupervised Domain Adaptation
Typically, UDA designed for major computer vision problems occupy three groups: domain discrepancy alignment, adversarial learning, and self-training. In domain discrepancy alignment, a proper divergence measure, such as maximum mean discrepancy [51], correlation alignment [55], or Wasserstein distance [37] is chosen as a measure to minimize domain discrepancy in a latent feature space. Adversarial learning approaches [23, 22, 57, 53, 18, 61] aims to achieve domain invariance in rather input space, feature extraction, or output space. Instead of reducing domain discrepancy with a divergence measure, self-training methods [24, 25, 56, 64, 26] iteratively learn a segmentation model by generating and assigning pseudo-labels to model training in the unlabeled target domain. DaFormer [24] proposes a transformer-based domain adaptation model, effectively handling multi-domain data by introducing domain attention and adversarial training mechanisms. HRDA [25] is a model designed for domain adaptation in high-resolution images, enhancing domain adaptation performance through the integration of hierarchical attention and multi-resolution representations. MIC [26] leverages masked image consistency to enhance context understanding for generalization. In this work, our proposed solution follows the pattern of self-training, assigning pseudo-labels to unlabeled training examples through a model learned by the labeled data, and iteratively including them in the learning process.
Specifically, the main purpose of this work is to explore UDA for instance segmentation in the field of autonomous driving. Previous work exploring UDA methods for image segmentation tasks mainly focuses on semantic segmentation [24, 25, 28]. Recently, some work on UDA for panoptic segmentation [27, 52] has been proposed that can simultaneously achieve domain-adaptive instance and semantic segmentation within a single network. Nevertheless, they are not designed for instance segmentation. The most related to our work is [70], which also aims to address the challenge of synth-to-real domain adaptation for instance segmentation. It consists of three feature adaptation modules at different levels (global, local, and subtle), and the training process was based on discriminator networks with adversarial learning. However, this work adopts the MaskR-CNN framework, whose performance might be limited by the CNN architecture and the R-CNN framework designed for the object detection task, thus leading to non-ideal results compared to transformer-based models. In contrast, we focus on designing a synth-to-real UDA framework following the self-training pattern with an advanced transformer-based segmentation network Mask2Former.
@@ 2.3 Cross-domain Data Mixing
A commonly used data augmentation technique for classification and semantic segmentation tasks is mixing [33]. There are two main types of mixing: soft mixing [69, 19, 13] and hard mixing [68, 59, 62, 56, 32, 25, 30]. The former one obtains mixed images by pixel-wise weighted average of two images, while the latter extracts patches from other image and pastes it to the current image. In UDA, this augmentation is usually applied in the form of cross-domain data mixing, i.e. the mixed samples consist of data from both source and target domains. Tranheden et al. introduce DACS [56], a unidirectional cross-domain data mixing strategy that cuts out semantic pixels in the source domain and pastes them into samples in the target domain. This strategy is further applied in subsequent research [25]. Gao et al. propose a dual soft-paste strategy that not only considers mixing data from the source to target images but also within the source domain itself [13]. A similar idea is suggested by Huo et al., while for the dual mixing[30], they focus on the target-to-target domain. Our proposed strategy differs from the aforementioned papers in applying a bidirectional domain mixing which includes both S2T and T2S directions. Our intuition is that incorporating T2S data mixing can bootstrap the model to better learn features in the target domain, thereby improving domain adaptation performance. Most recently, Kim et al. also propose a bidirectional cut-and-paste strategy based on image patches [32]. The main difference between our strategy and theirs is that we cut out instance objects instead of image patches. We believe that instance-based data mixing can reduce the impact of redundant background features on instance segmentation compared to patch-based ones.
@@ 3.1 Overall Approach
For the source domain dataset , comprising images and their instance labels , an instance segmentation model can be trained in a supervised manner. However, the target domain dataset only provides unlabeled images . We aim to address the domain gap by leveraging , , and to train a model capable of accurately predicting object instances on .
@@
then the network can be used to predict pseudo-labels on target domain . In our case, we use an advanced universal image segmentation architecture, Mask2Former[6], as the base network for instance segmentation. The segmentation loss for Mask2Former is
where binary cross-entropy loss and dice loss is for binary mask segmentation, is for instance class classification. , and are the hyper-parameters.
In the second stage, we utilize the cross-domain bidirectional mixed images and labels produced by both source and target domain data to resume the training of the model for iterations. The rare-class instances in the source domain are resampled during data mixing. Since the target images are unlabeled, we utilize the pseudo-labels generated by the exponential moving average (EMA) teacher model . At the beginning of the second stage, the parameters of the EMA teacher are initialized by the weights of . Then, in each training step, we first infer the target images by the EMA teacher to get pseudo-labels . In line with common practice [56, 9, 25, 24, 26], we use confidence thresholds to filter out noisy pseudo-labels. A pseudo-label confidence score is calculated as the product of mask confidence (confidence of whether an instance object exists in a mask) and class confidence (confidence of an instance class). Then, we leverage the ground-truth labels and the pseudo-labels to perform bidirectional instance cut-and-paste, thereby obtaining mixed data . Finally, mixed data are used to train the model with a weighted bidirectional data mixing loss
where and are weights for different directional mixing data. Then the EMA teacher parameters are updated. Algorithm 1 illustrates the flow of our proposed approach in detail.
@@
In data mixing, the mixed images are generated by cut-and-paste image regions from one image to another. Different from DACS [56] which only performs unidirectional data mixing in the S2T direction (see Fig. 5), we propose an instance-level bidirectional data mixing strategy that fully utilizes instances from both source and target domains (see Fig. 5). As shown in Fig. 5, in each iteration of the mixed data training stage, the mixed data is generated in both S2T and T2S directions. For S2T mixing, the instances in a source domain image are cut and pasted onto a target domain image , resulting in an augmented mixing images . Its corresponding instance label is created according to the image, where the instance labels are added to the generated target pseudo-labels . Similarly, to obtain the T2S mixing data and , we leverage the generated pseudo-labels of the target images to mask out the instances that are desired to be pasted on the source image . The network is trained on mixed images , with mixed labels , . Compared to using image patches for mixing, we use instance-based data mixing to avoid the impact of redundant background, leading to a more realistic mixed image. We provide rich experimental results in Sec. 4 to demonstrate the effectiveness of our proposed bidirectional data mixing strategy.
[2]
[…]
{subfloatrow}[1]
[…]
@@
Conventionally, the instance class labels have specific semantic meanings such as: ‘Motorcycle’, ‘Bicycle’, ‘Train’ etc. Instances with similar semantics tend to have more common features. For example, ‘Motorcycle - Bicycle’ may contain more common features than ‘Train - Bicycle’, therefore leading to a higher inter-class correlation. Based on this inductive bias, we propose category module training. Instead of training one model to predict all classes, we separate the classes into several category groups. We train one model for each group and combine the outputs of these models to obtain the final result. Our assumption is that introducing this inductive bias may guide the model to better learn and associate the inter-class and intra-class features. We provide rich ablation study results in Sec. 4 to verify the idea. Note that in this work, we consider the number of category groups to two, to verify the profit of the proposed category module training.
@@
Class imbalance has a negative impact on the performance of deep learning models. Since the goal of model training is to minimize the overall error, models trained on datasets with obvious class imbalance tend to be biased toward the classes with large amounts of data, resulting in poor performance on the classes with few data. In our three synthetic datasets, we observe that there exists a class imbalance problem. Therefore, we want to increase the training frequency of instances from these rare classes by intentionally cutting and pasting them in data mixing. In each category group, for the class that occupies the smallest percentage of instances, we define it as a rare class: ‘Motorcycle’ and ‘Train’ for UrbanSyn (2.47% and 0.37%) and Synscapes (5.91% and 2.55%), ‘Motorcycle’ and ‘Bus’ for SYNTHIA (2.18% and 2.08%).
Fig. 6 demonstrates the operation of our proposed rare-class balancing. Specifically, during the process of traversing images for data mixing, for those images containing rare-class instances, we store these images and the corresponding instance labels in a pool with a maximum number of sample N. These N samples in the pool are continuously updated with newly emerged rare-class instance objects and their corresponding images in a FIFO (first-in, first-out) fashion. In turn, when encountering images that do not contain rare classes, we randomly select half of the samples in the pool and mix them on these images. To avoid error accumulation from pseudo-labels, we only apply rare-class balancing to data mixing in the S2T manner. Note that the whole process is done online, thus no pre-processing is required.

### v3

# UDA4Inst: Unsupervised Domain Adaptation for Instance Segmentation

@@ Abstract
Unsupervised Domain Adaptation (UDA) aims to transfer knowledge learned from a labeled source domain to an unlabeled target domain. While UDA methods for synthetic to real-world domains (synth-to-real) show remarkable performance in tasks such as semantic segmentation and object detection, very few were proposed for instance segmentation in the field of vision-based autonomous driving, and the existing ones are based on a suboptimal baseline, which severely limits the performance. In this paper, we introduce UDA4Inst, a strong baseline of synth-to-real UDA for instance segmentation. UDA4Inst adopts cross-domain bidirectional data mixing at the instance level to effectively utilize data from both source and target domains. Rare-class balancing and category module training are also employed to further improve the performance. It is worth noting that we are the first to demonstrate results on two new synth-to-real instance segmentation benchmarks, with 39.0 mAP on UrbanSyn Cityscapes and 35.7 mAP on Synscapes Cityscapes. Our method outperforms the source-only Mask2Former model by +7 mAP and +7.6 mAP, respectively. On SYNTHIA Cityscapes, our method improves the source-only Mask2Former by +6.7 mAP, achieving state-of-the-art results.111Our code will be released soon.
@@
Synth-to-Real Unsupervised Domain Adaptation Instance Segmentation Vision-based Autonomous Driving.1 Introduction
@@
Instance segmentation provides a detailed and fine-grained understanding of the objects in an image, which is critical for a variety of real-world applications, such as multi-object tracking [24, 41, 2], pose estimation [6, 1], and motion prediction [51, 29]. Specifically, in autonomous driving, instance segmentation is useful as it helps to locate safety-critical objects (such as pedestrians or riders) with accurate boundaries, enabling more advanced scene analysis and decision-making in many scenarios [9, 49, 2, 52].
@@
However, in practice, training instance segmentation models requires access to a large amount of densely labeled data. Each instance needs to be accurately delineated with pixel-level masks, which is laborious and highly time-consuming. In addition, to ensure proper generalization of the model, the data should contain a wide range of object appearances, poses, scales, and backgrounds. Therefore, the label acquisition and object variability needed for instance segmentation encounter a data-hungry bottleneck. This challenge demands a more elaborate approach to enhance the efficiency of the labeling of instance segmentation data.
@@
As an alternative, synthetic data are increasingly being used to train deep models as they enable large-scale training data, which is especially relevant for vision tasks [38]. Synthetic data offers labels annotated directly by simulators for free [10, 37]. Crucially, it allows for the incorporation of diversity into the training dataset, thereby enhancing the model’s generalization capability.
@@
However, when training models with synthetic data and testing them in the real-world domain, the results are often unsatisfactory due to the inevitable gap between the two domains. UDA methods are essential for mitigating this inherent gap between the synthetic source and the real-world target domains while avoiding human labeling. In recent years, many studies have explored the integration of synthetic data and UDA methods to address a wide range of computer vision tasks including image classification [11, 33, 40, 35], object detection [46, 13, 15, 39, 54, 36], and semantic segmentation [3, 19, 20, 23, 16, 17, 32]. Comparatively, there has been relatively little research on UDA for instance segmentation, reflected in only three works [55, 22, 44] in the last five years. Among them, although [22, 44] do provide results on instance segmentation, they were models designed for the panoptic segmentation task, and [55] was based on a suboptimal Mask R-CNN framework with adversarial learning.
@@
In this paper, we focus on designing a framework that provides a strong baseline on UDA for Instance Segmentation (UDA4Inst) following an online self-training pattern. Unlike previous models [11, 20, 24] that perform typical unidirectional cross-domain data mixing, i.e., either source-to-target (S2T) or target-to-source (T2S), UDA4Inst utilizes an instance-based bidirectional cross-domain data mixing approach tailored for instance segmentation. For the T2S phase, we need to take pseudo labels for real-world instances and accommodate them into synthetic images. When multiple instances of the same class overlap, in semantic segmentation, these overlapping instances are treated as a single blob. In contrast, instance segmentation methods need to find clear boundaries in the overlapping region of objects to separate them. Note that this is especially more difficult than semantic segmentation. Additionally, to mitigate the problem of instance imbalance in existing datasets, UDA4Inst incorporates a handy online algorithm to balance rare-class instances during data mixing between domains. In this way, we observe a clear improvement in the segmentation performance of rare-class instances. As the third methodological contribution, we explore category module training, where multiple classes are categorized into several modules for separate training, then we concatenate outputs of these modules to obtain a final prediction. We demonstrate the effectiveness of UDA4Inst on three UDA instance segmentation benchmarks. To the best of our knowledge, we are the first to perform UDA instance segmentation using the synthetic driving scene datasets UrbanSyn [14] and Synscapes [53], achieving 39.0 mAP and 35.7 mAP on Cityscapes [9]. Additionally, we have achieved state-of-the-art results on SYNTHIA [42] Cityscapes. Compared to the source-only method, our approach shows improvements of +7 mAP, +7.6 mAP, and +6.7 mAP on these three benchmarks.
@@ 2.1 Instance Segmentation
Early mainstream deep learning models based on convolutional neural networks (CNNs) for instance segmentation, such as Mask R-CNN [18] and YOLACT [4], have achieved remarkable successes. In recent years, with the adoption of the self-attention mechanism in natural language processing and computer vision, transformer-based models have emerged and surpassed CNN-based architectures in many tasks such as object detection [28], semantic segmentation [8, 7], and panoptic segmentation. [44, 7]. Among them, Mask2Former [7] has attracted great attention as a universal transformer-based architecture that obtains top results in three major image segmentation tasks (panoptic, instance, and semantic) on several benchmarks. It introduces masked attention, which extracts localized features by constraining cross-attention within predicted mask regions. In this work, we leverage this advanced universal segmentation architecture to prompt the state-of-the-art results in instance segmentation with UDA.
@@ 2.2 Unsupervised Domain Adaptation
Typically, UDA methods occupy three groups: domain discrepancy alignment, adversarial learning, and self-training. In domain discrepancy alignment, a proper divergence measure, such as maximum mean discrepancy [43], correlation alignment [47], or Wasserstein distance [31] is chosen to minimize domain discrepancy in a latent feature space. Adversarial learning approaches [45, 50] aim to achieve domain invariance in input space, feature extraction, or output space. To reduce domain discrepancy, self-training methods [19, 20, 21] train a model using a combination of labeled and unlabeled data. These methods generate high-confidence pseudo-labels from unlabeled data and iteratively incorporate them to expand the labeled data during training, gradually improving the model’s performance. HRDA [20] enhances domain adaptation performance through the integration of hierarchical attention and multi-resolution representations. MIC [21] leverages masked image consistency to enhance context understanding for generalization. Our proposed solution follows the pattern of self-training since it is the mainstream approach in state-of-the-art UDA methods, and is most suitable for our bidirectional mix strategy.
Current UDA works [19, 20, 23, 16, 17, 32] mainly focus in the semantic segmentation task that is getting really close to the supervised learning methods trained on human labeling datasets. On the other hand, UDA panoptic segmentation is yet to be thoroughly explored, where few works [22, 44] can simultaneously achieve domain-adaptive instance and semantic segmentation within a single network. The work most related to ours is [55], which proposes a multi-level feature adaptation strategy to improve synth-to-real instance segmentation. To reduce the domain gap, they focus on aligning feature distributions between synthetic and real data at global, local, and subtle levels. This is achieved by using Mask R-CNN and discriminator networks with adversarial learning. In recent years, there has been significant progress in the research fields of non-cross-domain instance segmentation and UDA methods. In this work, we aim to improve the baseline on UDA for instance segmentation with approaches such as cross-domain data mixing along with the advanced transformer-based Mask2Former network [7].
@@ 2.3 Cross-domain Data Mixing
A commonly used data augmentation technique is mixing image content. In UDA, this augmentation is usually applied in the form of cross-domain mixing, i.e. the mixed samples consist of data from both source and target domains. Tranheden et al. introduce DACS [48], a unidirectional cross-domain data mixing strategy that cuts out semantic classes in the source domain and pastes them into samples in the target domain. Suhyeon et al. use a similar method by transferring the tail-class content from source to target to solve the class imbalance problem [27]. Gao et al. propose a dual soft-paste strategy that not only considers mixing from the source to target images but also within the source domain itself [12]. A similar idea is suggested by Huo et al., while for the dual mixing [25], they focus on the target-to-target domain. Most recently, Kim et al. propose a S2T and T2S cut-and-paste strategy based on predefined image patches [26]. It is noteworthy that these mixing approaches fall into two categories: 1) Class-wise Mixing. For instance, Cardace et al. [5] only mix semantic classes such as ‘Car’, ‘Person’, and ‘Pole’. This approach fails to delineate boundaries accurately where multiple instances overlap. 2) Random Patches Mixing which involves cut-and-paste based on arbitrary patches, resulting in unrealistic images with mixed foreground instances and background area.
Unlike the prior works in semantic segmentation, we propose a bidirectional domain mixing strategy tailored for instance segmentation, which involves both S2T and T2S mixing. Since our work focuses on instance segmentation, we perform cut-and-paste based on semantic instances rather than classes or random patches. Therefore, instances pasted in the new domain do not include background pixels from the original domain, generating more realistic mixed data.
@@ 3.1 Overall Approach
For the source domain dataset , comprising images and their instance labels , an instance segmentation model can be trained in a supervised manner. However, the target domain dataset only provides unlabeled images . We aim to address the domain gap by leveraging , , and to train a model capable of accurately segmenting instances on .
@@
then the network can be used to predict pseudo-labels, , on the target domain. In our case, we use Mask2Former [7] as the base network for instance segmentation. The segmentation loss for Mask2Former is
where binary cross-entropy loss and dice loss are for binary mask segmentation, is for instance class classification. , and are weighting hyper-parameters.
In the second stage, we utilize the cross-domain bidirectional mixed images and labels produced by both source and target domain data to resume the training of the model for iterations. The rare-class synthetic instances defined in the source domain are resampled during data mixing. Since the target images are unlabeled, we utilize the pseudo-labels generated by the exponential moving average (EMA) teacher model . At the beginning of the second stage, the parameters of the EMA teacher are initialized by the weights of . Then, in each training step, we first infer the target images by the EMA teacher to get pseudo-labels . In line with common practice [48, 20, 19, 21], we use confidence thresholds to filter out noisy pseudo-labels. A pseudo-label confidence score is calculated as the product of mask confidence (confidence of whether an instance object exists in a mask) and class confidence (confidence of an instance class). Then, we leverage the ground-truth labels and the pseudo-labels to perform bidirectional instance cut-and-paste, thereby obtaining mixed data . Finally, mixed data are used to train the model with a weighted bidirectional data mixing loss
where and are weights for different directional mixing data. Then the EMA teacher parameters are updated. We present the algorithm process in the supplementary material.
@@
In data mixing, the mixed images are generated by cut-and-paste image regions from one image to another. We adopt a bidirectional mixing strategy that fully utilizes instances from both source and target domains to facilitate instance segmentation. As shown in Fig. 2, in each iteration of the mixed data training stage, the mixed data is generated in both S2T and T2S directions. For S2T mixing, the instances in a source domain image are cut and pasted onto a target domain image , resulting in an augmented mixing images . Its corresponding instance label is created according to the image, where the instance labels are added to the generated target pseudo-labels . Similarly, to obtain the T2S mixing data and , we leverage the generated pseudo-labels of the target images to mask out the instances that are desired to be pasted on the source image . The network is trained on mixed images , with mixed labels , .
@@
Conventionally, the instance labels have specific semantic meanings such as: ‘Motorcycle’, ‘Bicycle’, ‘Train’ etc. Instances with similar semantics tend to have more common features. For example, ‘Motorcycle - Bicycle’ may contain more common features than ‘Train - Bicycle’, therefore leading to a higher inter-class correlation. Based on this inductive bias, we propose category module training. Instead of training one model to predict all classes, we separate the classes into several category groups. We train one model for each group and combine the outputs of these models to obtain the final result. We assume that introducing this inductive bias may guide the model to better learn and associate the inter-class and intra-class features. To verify the idea, we provide rich ablation study results in Sec. 4.2.4. In this work, we consider the number of category groups as two.
The category module training can greatly improve the accuracy, yet there is a trade-off between model performance and testing efficiency, as it requires two category modules for inference. Considering the testing efficiency, we also provide an alternative that only uses one model during testing, which can achieve comparable performance. In this way, we use the two-category models to generate pseudo-labels during the self-training phase and train a single model which is then used for testing. We name this method as category-free and provide results in the Sec. 4.2.3.
@@
Class imbalance is a common issue that harms the performance of deep learning models. Models trained on imbalanced datasets tend to favor classes with more data, leading to poor performance in classes with fewer examples. This problem exists in the three synthetic datasets we used. Therefore, we increase the training frequency of these rare class instances by intentionally resampling them in data mixing. In each category group, for the class that occupies the smallest percentage of instances in the source dataset, we define it as a rare class: ‘Motorcycle’ and ‘Train’ for UrbanSyn (2.47% and 0.37%) and Synscapes (5.91% and 2.55%), ‘Motorcycle’ and ‘Bus’ for SYNTHIA (2.18% and 2.08%).
Specifically, during the process of traversing images for data mixing, for those images containing rare-class instances, we store these images and the corresponding instance labels in a pool with a maximum number of N. These samples are continuously updated with newly emerged rare-class instances and their corresponding images in a FIFO (first-in, first-out) fashion. In turn, when encountering images that do not contain rare classes, we randomly select half of the samples in the pool and mix them on these images. To avoid error accumulation from pseudo-labels, we only apply rare-class balancing to data mixing in the S2T direction. Note that unlike [21], we do not need to pre-process data to gather information about the distribution of instances for each class in the dataset.
| AP | AP50 | |||||||||||||||||||
| Model | Person | Rider | Car | Truck | Bus | Train | M.C | B.C | mAP | Imp. | Person | Rider | Car | Truck | Bus | Train | M.C | B.C | mAP50 | Imp. |
| upper bound | 41.1 | 32.4 | 61.4 | 44.8 | 67.5 | 50.7 | 26.8 | 26.6 | 43.9 | - | 75.4 | 71.3 | 87.2 | 58.7 | 80.3 | 76.1 | 58.1 | 65.1 | 71.5 | - |
| UrbanSyn Cityscapes | ||||||||||||||||||||
| w/o UDA | 25.1 | 19.5 | 39.8 | 37.6 | 51.5 | 46.7 | 19.2 | 16.2 | 32.0 | - | 52.0 | 48.4 | 61.5 | 48.6 | 66.2 | 70.0 | 42.6 | 46.0 | 54.4 | - |
| UDA4Inst | 30.0 | 26.4 | 49.2 | 42.9 | 64.4 | 58.1 | 21.3 | 19.7 | 39.0 | +7.0 | 60.1 | 60.7 | 75.1 | 56.4 | 78.0 | 77.8 | 48.7 | 53.6 | 63.8 | +9.4 |
| Synscapes Cityscapes | ||||||||||||||||||||
| w/o UDA | 27.2 | 19.2 | 39.5 | 24.0 | 46.7 | 33.1 | 18.7 | 16.6 | 28.1 | - | 53.9 | 44.3 | 67.8 | 31.9 | 60.6 | 58.1 | 39.9 | 45.4 | 50.2 | - |
| UDA4Inst | 30.3 | 27.1 | 47.5 | 34.5 | 58.5 | 46.3 | 22.3 | 19.1 | 35.7 | +7.6 | 59.9 | 63.2 | 77.9 | 45.3 | 72.8 | 63.5 | 46.1 | 51.7 | 60.1 | +9.9 |
| SYNTHIA Cityscapes | ||||||||||||||||||||
| Zhang [55] w/o UDA | 7.5 | 0.6 | 19.1 | - | 17.0 | - | 2.5 | 0.7 | 7.9 | - | 25.2 | 4.3 | 39.3 | - | 31.3 | - | 9.6 | 4.0 | 18.9 | - |
| Zhang [55] | 9.4 | 3.8 | 22.2 | - | 23.5 | - | 3.2 | 2.4 | 10.8 | +2.9 | 28.8 | 17.8 | 39.9 | - | 37.5 | - | 12.8 | 11.9 | 24.8 | +5.9 |
| CVRN [22] w/o UDA | - | - | - | - | - | - | - | - | - | - | 30.4 | 9.2 | 32.5 | - | 20.4 | - | 3.1 | 1.3 | 16.2 | - |
| CVRN [22] | - | - | - | - | - | - | - | - | - | - | 34.4 | 25.3 | 38.7 | - | 38.1 | - | 10.1 | 8.7 | 25.9 | +9.7 |
| EDAPS [44] w/o UDA | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 22.0 | - |
| EDAPS [44] | 23.5 | 13.4 | 27.5 | - | 23.6 | - | 5.8 | 0.1 | 15.7 | - | 53.9 | 40.6 | 50.5 | - | 38.8 | - | 21.1 | 0.7 | 34.3 | +12.3 |
| w/o UDA | 26.3 | 15.2 | 38.3 | - | 43.0 | - | 13.6 | 10.9 | 24.6 | - | 53.4 | 40.8 | 68.1 | - | 56.5 | - | 36.4 | 36.1 | 48.6 | - |
| UDA4Inst | 28.1 | 19.6 | 47.9 | - | 58.7 | - | 18.5 | 15.0 | 31.3 | +6.7 | 57.7 | 52.5 | 76.9 | - | 68.6 | - | 42.6 | 43.3 | 56.9 | +8.3 |